OpenStudy (anonymous):

In a single-slit diffraction pattern, the distance between the first and second minima is 1.8cm on a screen 4.4m from the slit. Find the slit width given that the wavelength of the light is 550nm.

OpenStudy (ash2326):

@Mythias it won't help much posting this here. Please post this in Physics. Thanks:)

OpenStudy (anonymous):

its just that the physics forum is dead....lol plus ive always gotten more help here LOL

OpenStudy (callisto):

First, you should put this question in the Physics group. Second, slit width = λD/a , if you remember it correctly

OpenStudy (callisto):

sorry , you -> I :S

OpenStudy (experimentx):

seems like double slit interference

OpenStudy (callisto):

that would be dsinθ = mλ, i believe

OpenStudy (anonymous):

ya the formulas u guys have r correct

OpenStudy (experimentx):

that looks like bragg's law ... if you add 2 on lhs ...

OpenStudy (anonymous):

do we do y=mlamdaL/d?

OpenStudy (callisto):

sorry i remember that wrong :S fringe separation = λD/a, a is the slit separation, d is the distance from the screen ...

OpenStudy (callisto):

1.8/100 = ( 550x10^-9 ) (4.4) /a a = ( 550x10^-9 ) (4.4) / 0.018

OpenStudy (experimentx):

yeah you are right .. same as iterference comes form here http://hyperphysics.phy-astr.gsu.edu/hbase/phyopt/imgpho/ymin2.gif

OpenStudy (callisto):

I don't like doing these king of question, i mixed the formulas up or can't remember the unknowns well :S

OpenStudy (callisto):

*kind

OpenStudy (anonymous):

thanks guys :)